Autumn:

How many moles of water are in 8.56x10 24 atoms of water

sillybilly123:

this qu .... you might want to post a screengrab? \(1 \text{mol}_\text{water} \equiv 6 \times 10^{23} \text{ molecules of water } \equiv 18 \times 10^{23} \text{ atoms}\) So: \(8.56 \times 10^{ 24} \text{atoms} \equiv \dfrac{8.56 \times 10^{ 24}}{18 \times 10^{23}} \text{mol} \approx 4.76 \text{ mol}\)

Vocaloid:

this is kind of a weirdly phrased question since we wouldn't normally call them "atoms of water" it would be "molecules of water"